OpenStudy (anonymous):

How can I solve for \(f'(x)\) if \(f(x) = tan(\frac{\pi x}{2})\) and \(0

OpenStudy (anonymous):

chain rule

OpenStudy (anonymous):

I got \(f'(x) = \frac{cos^2(\frac{\pi x}{2})\frac{\pi}{2}+sin^2(\frac{\pi x}{2})}{cos^2(\frac{\pi x}{2})}\)

OpenStudy (anonymous):

how can i simplify that?

OpenStudy (anonymous):

well...how u get that? \[(\tan u)'=u' (1+\tan^2 u)\]

OpenStudy (anonymous):

tanx = sinx/cosx

OpenStudy (anonymous):

oh yeah ... there is a little typo there\[f'(x) =\frac{\pi}{2} \frac{\cos^2(\frac{\pi x}{2})+\sin^2(\frac{\pi x}{2})}{\cos^2(\frac{\pi x}{2})}\]

OpenStudy (anonymous):

and for simplifying\[\sin^2 \text{everything}+\cos^2 \text{everything}=1\]:)

OpenStudy (anonymous):

Looks like i made a mistake there... the next part of the question is to find the equation of the tangent line to this graph at x = 1/3. How do I take the domain into consideration?

OpenStudy (anonymous):

well we have\[f'(x) =\frac{\pi}{2} \frac{1}{\cos^2(\frac{\pi x}{2})}\] right?

OpenStudy (anonymous):

x=1/3 is in our domain and no problem with workin on that

OpenStudy (anonymous):

\(f'(\frac{1}{3}) = \frac{\pi}{2cos^2(\frac{\pi}{6})}\)

OpenStudy (anonymous):

well thats the tangent line slope

OpenStudy (anonymous):

how can i find the other variables of y = mx+c?

OpenStudy (anonymous):

the line is tangent to the \(f(x)\) so the point \((1/3,f(1/3))\) lies on the line also

OpenStudy (anonymous):

line intersects \(f(x)\) at the point \(x=1/3\) a i right?

OpenStudy (anonymous):

*am

OpenStudy (anonymous):

I'm confused, how can I find the y intercept of f'(x) from here?

sam (.sam.):

Just apply chain rule here Use \[\frac{ d }{ dx }(\tan(x))=\sec^2(x)\] -------------------------------------------------------------------- \[\frac{d}{dx}\left(\tan \left(\frac{\pi x}{2}\right)\right)=\frac{1}{2} \pi \sec ^2\left(\frac{\pi x}{2}\right)\]

OpenStudy (anonymous):

I can't use secant in this question though

sam (.sam.):

You mean you want to find the angles?

OpenStudy (anonymous):

just find the equation of the tangent

OpenStudy (anonymous):

y=mx+c we know the value of m right till here?

OpenStudy (anonymous):

yes, from the derivative

OpenStudy (anonymous):

ok u just need to evalute c ... if we obtain a point of line we can find c right?

OpenStudy (anonymous):